# Black-Box On-Policy Distillation of Large Language Models

### v1

@@ 2.2 Training
We find that jointly warming up the generator and discriminator before the GAD training stage is crucial for final performance. We fine-tune the student on the teacher’s response, and we minimize the cross-entropy loss as warmup for the generator. In the meanwhile, the discriminator is trained using the same data with the Bradley-Terry loss in Equation (3). We conduct warmup for both models for one epoch before starting GAD training. This step promotes effective adversarial optimization and ensures the balance between the generator and discriminator. Ablation studies on the warmup strategy are presented in Section˜3.3.
@@ 3.3 Analysis
We perform an ablation study of the warmup strategy introduced in Section˜2.2. As shown in Table˜3, we separately remove the warmup stage for the generator and the discriminator on Qwen2.5-7B-Instruct. When removing the generator warmup, we directly use Qwen2.5-7B-Instruct without SeqKD as initialization for both the generator and discriminator for GAD training. This leads to a performance drop. We attribute this to the discriminator easily distinguishing between the student and teacher outputs in the early training stage. The large distributional gap between the teacher and the student weakens the effectiveness of GAD training. When removing the discriminator warmup, we use the generator obtained after one epoch of SeqKD and initialize the discriminator with the original Qwen2.5-7B-Instruct. In this setting, the imbalance between the generator and the discriminator prevents the discriminator from providing sufficiently informative feedback. Consequently, the adversarial interaction becomes ineffective, and the generator exhibits little improvement beyond its warmup performance.
@@ A.2 Training Details
We search learning rate in [1e-6, 5e-6] for GAD and SeqKD baseline. For SeqKD, we find 5e-6 leads to better results in all experiments. For GAD with GPT-5-Chat teacher, we use 1e-6 for both warmup and GAD training stage, and for GAD with Qwen2.5 teacher as in Table˜5, we use 5e-6 for warmup stage and 1e-6 for GAD training stage. The maximum context length is set to 2048 tokens for instruction prompts and 1536 for model responses. The training temperature is set to $0.8$.
@@ B.1 Additional Automatic Evaluation Results
We provide additional results of the automatic evaluation. In Table˜4, we report GPT-4o score and response lengths of distilled student models trained with the GPT-5-Chat teacher. Across datasets, we observe that SeqKD tends to produce shorter responses that closely follow the teacher’s length distribution whereas GAD maintains the original model’s length distribution while integrating the teacher’s global stylistic characteristics. We attribute this behavior to the on-policy sampling of GAD, which encourages generation patterns aligned with both the student’s prior and the teacher’s guidance.
@@
In Table˜5, we distill from Qwen2.5-14B-Instruct teacher to student models from the Llama family. Although the teacher is open-source, its tokenizer is incompatible with the students, preventing direct application of white-box distillation methods that align KL divergence between teacher and student logits. In this setting, GAD remains effective, outperforming both the pre-distillation models and the SeqKD baseline in most settings on GPT-4o evaluation score.

### v2

@@ 2.2 Training
We find that jointly warming up the generator and discriminator before the GAD training stage is crucial for final performance. We fine-tune the student on the teacher’s response, and we minimize the cross-entropy loss as warmup for the generator. In the meanwhile, the discriminator is trained using the same data with the Bradley-Terry loss in Equation (3). We conduct warmup for both models for one epoch before starting GAD training. This step promotes effective adversarial optimization and ensures the balance between the generator and discriminator. Ablation studies on the warmup strategy are presented in Section˜3.4.
@@ 3.4 Ablations
LMSYS Others
SeqKD 47.5 46.2
GAD
   Disc. BT Loss (Default) 48.9 47.9
   Disc. CE Loss 47.9 46.4
Table 4: Ablation of discriminator loss choice on Qwen2.5-3B-Instruct. Default Bradley-Terry loss outperforms cross-entropy loss.
@@
We perform an ablation study of the warmup strategy introduced in Section˜2.2. As shown in Section˜3.4, we separately remove the warmup stage for the generator and the discriminator on Qwen2.5-7B-Instruct. When removing the generator warmup, we directly use Qwen2.5-7B-Instruct without SeqKD as initialization for both the generator and discriminator for GAD training. This leads to a performance drop. We attribute this to the discriminator easily distinguishing between the student and teacher outputs in the early training stage. The large distributional gap between the teacher and the student weakens the effectiveness of GAD training. When removing the discriminator warmup, we use the generator obtained after one epoch of SeqKD and initialize the discriminator with the original Qwen2.5-7B-Instruct. In this setting, the imbalance between the generator and the discriminator prevents the discriminator from providing sufficiently informative feedback. Consequently, the adversarial interaction becomes ineffective, and the generator exhibits little improvement beyond its warmup performance.
Discriminator Loss Choice
We ablate the choice of discriminator loss in Section˜3.4, and observe that our default Bradley-Terry loss outperforms cross-entropy loss in overall GPT-4o evaluation score. The Bradley-Terry loss is defined in Equation (3), while the cross-entropy loss is a binary classification loss commonly adopted for discriminator in prior works [gan, seqgan, GAIL, learningdense]. The cross-entropy discriminator loss can be written as:
$\min_{D}\ \mathbb{E}_{(x,y_{t})\sim\mathcal{T}}\left[-\log\sigma\left(D(y_{t})\right)-\log\left(1-\sigma\left(D(G(x))\right)\right)\right].$ (4)
Experiments on Qwen2.5-3B-Instruct shows that Bradley-Terry loss can enhance discriminator training stability and improve automatic evaluation scores over cross-entropy loss. The result highlights the effectiveness of Bradley-Terry loss for discriminator training in LLMs.
Gen. Size Disc. Size LMSYS Others
3B 3B (Default) 48.9 47.9
3B 7B 47.8 46.9
7B 7B (Default) 50.8 50.0
7B 14B 50.5 49.9
Table 5: Ablation of discriminator model size. Experiments are performed on Qwen2.5 Instruct models, evaluated with GPT-4o score.
Discriminator Model Size
We ablate the relative model size of the generator and discriminator in Table˜5. Using equal model sizes for the two models, which is our default setting, yields the best performance. The experiments are conducted on Qwen2.5 Instruct models and evaluated by GPT-4o scores. We increase the discriminator model size from 3B to 7B for the 3B student, and from 7B to 14B for the 7B student in GAD. Increasing the discriminator size does not improve performance. The experiment shows maintaining a balanced generator-discriminator pair is crucial for achieving strong performance.
@@ A.2 Training Details
We search learning rate in [1e-6, 5e-6] for GAD and SeqKD baseline. For SeqKD, we find 5e-6 leads to better results in all experiments. For GAD with GPT-5-Chat teacher, we use 1e-6 for both warmup and GAD training stage, and for GAD with Qwen2.5 teacher as in Table˜7, we use 5e-6 for warmup stage and 1e-6 for GAD training stage. The maximum context length is set to 2048 tokens for instruction prompts and 1536 for model responses. The training temperature is set to $0.8$.
@@ B.1 Additional Automatic Evaluation Results
We provide additional results of the automatic evaluation. In Table˜6, we report GPT-4o score and response lengths of distilled student models trained with the GPT-5-Chat teacher. Across datasets, we observe that SeqKD tends to produce shorter responses that closely follow the teacher’s length distribution whereas GAD maintains the original model’s length distribution while integrating the teacher’s global stylistic characteristics. We attribute this behavior to the on-policy sampling of GAD, which encourages generation patterns aligned with both the student’s prior and the teacher’s guidance.
@@
In Table˜7, we distill from Qwen2.5-14B-Instruct teacher to student models from the Llama family. Although the teacher is open-source, its tokenizer is incompatible with the students, preventing direct application of white-box distillation methods that align KL divergence between teacher and student logits. In this setting, GAD remains effective, outperforming both the pre-distillation models and the SeqKD baseline in most settings on GPT-4o evaluation score.